A 10ft long horizontal through has ends in the shape of equilateral triangles. The sides of the triangle are 6ft in length. If water is being poured inside the through, find a mathematical model that expresses the volume of the water as a function of its depth


1
Expert's answer
2021-10-12T01:53:49-0400

Volume=cross sectional area*length

C.S.A="\frac{1}{2}bh=\frac{1}{2}b*bSin60°"

With l as length,

Volume,V="\frac{1}{2}b*bSin60°*l"

Mathematical model will have a partial derivative with respect to l

"\frac{\partial l}{\partial b}(V)"

="\frac{\partial l}{\partial b}(\frac{1}{2}b*bSin60°*l)"

"=\frac{1}{4}b*bSin60°*l^{2}"